# Ca2-VDM: Efficient Autoregressive Video Diffusion Model with Causal Generation and Cache Sharing

### v1

@@ Appendix C Training Details and Hyperparameters
Text-to-Video (T2V) Training. We trained Ca2-VDM and the OS-Fix baseline on a large-scale video-text dataset InternVid [45], by filtering it to a sub-set of 4.9M high-quality video-text pairs with resolution 256$\times$256. For Ca2-VDM, the training consists of two stages. We first train the causal modeling ability without the clean prefix (i.e., without conditional frames) on 32-frame videos. Then we use longer videos of 65 frames to train the model with the clean prefix, i.e., with $l=16$, $P_{\text{max}}=1+3l=49$ and $\max(L_{\text{train}})=P_{\text{max}}+l=65$. In the first stage, the model was trained with a batch size of 288 for 32k steps. In the second stage, it was trained with a batch size of 144 for 21k steps. For OS-Fix, it was trained with $L_{\text{train}}=32$ frames and $P=l=L_{\text{train}}/2=16$ frames, i.e., the prefix length is fixed. It was trained with a batch size of 288 for 20k steps  [^4] OS-Fix converges faster than Ca2-VDM since it only needs to learn fixed-length conditional frames..
Video Prediction Training. We trained OS-Fix, OS-Ext, and Ca2-VDM on the SkyTimelapse [51] dataset at resolution $256\times 256$ with $l=8$. OS-Ext and Ca2-VDM both used $P_{\text{max}}=1+3l=25$ (i.e., $L_{\text{train}}=33$). OS-Fix used a fixed $P=8$ and $L_{\text{train}}=16$. All three models were trained with a batch size of 8 for 11k steps  [^5] In contrast to text-to-video, the video prediction task on the SkyTimelapse dataset has less diversity and converges faster. So we used smaller batch size and training steps..
Hyperparameters. For all the training, we used the DDPM [16] schedule with $T=1000$, $\beta_{1}=10^{-4}$, and $\beta_{T}=0.02$. The models were trained using AdamW [20] optimizer with a learning rate of 2e-5. At the inference stage, we used the improved DDPM schedule [27] with 100 steps. For text-to-video, we set the classifier-free guidance scale as 7.5.
@@ D.1 Datasets
MSR-VTT [47]. we used its official test split which contains 2990 videos, with 20 manually annotated captions for each video. Following prior works [32, 50] and for fair comparisons, we randomly selected a caption for each video and generated 2990 videos for evaluation.
UCF101 [39]. As it only contains label names, we employed the descriptive text prompts from PYoCo [9], and generated 2048 samples with uniform distribution for each category following [14, 9, 32].
SkyTimelapse [51]. It is a time-lapse dataset showing dynamic sky scenes (e.g., cloudy sky with moving clouds). We used it for video prediction (i.e., without text input). Its training set contains 997 long timelapse videos, which are cut into 2392 short videos. Its test set contains 111 long timelapse videos, which are cut into 225 short videos. We trained the models on its training set and evaluated them on its test set.
@@ D.2 Quantitative Evaluation
Fréchet Video Distance (FVD) [40] measures the similarity between generated and real videos based on the distributions on the feature space. We followed prior works [3, 8, 32] to use a pretained I3D model [^6] https://github.com/songweige/TATS/blob/main/tats/fvd/i3d_pretrained_400.pt to extract the features. We used the codebase [^7] https://github.com/universome/stylegan-v from StyleGAN-V [35] to compute FVD statistics.

### v2

@@ Appendix C Training Details and Hyperparameters
Text-to-Video (T2V) Training. We trained Ca2-VDM and the OS-Fix baseline on a large-scale video-text dataset InternVid (Wang et al., 2024), by filtering it to a sub-set of 4.9M high-quality video-text pairs with resolution 256$\times$256. For Ca2-VDM, the training consists of two stages. We first train the causal modeling ability without the clean prefix (i.e., without conditional frames) on 32-frame videos. Then we use longer videos of 65 frames to train the model with the clean prefix, i.e., with $l=16$, $P_{\text{max}}=1+3l=49$ and $\max(L_{\text{train}})=P_{\text{max}}+l=65$. In the first stage, the model was trained with a batch size of 288 for 32k steps. In the second stage, it was trained with a batch size of 144 for 21k steps. For OS-Fix, it was trained with $L_{\text{train}}=32$ frames and $P=l=L_{\text{train}}/2=16$ frames, i.e., the prefix length is fixed. It was trained with a batch size of 288 for 20k steps  [^3] OS-Fix converges faster than Ca2-VDM since it only needs to learn fixed-length conditional frames..
Video Prediction Training. We trained OS-Fix, OS-Ext, and Ca2-VDM on the SkyTimelapse (Zhang et al., 2020) dataset at resolution $256\times 256$ with $l=8$. OS-Ext and Ca2-VDM both used $P_{\text{max}}=1+3l=25$ (i.e., $L_{\text{train}}=33$). OS-Fix used a fixed $P=8$ and $L_{\text{train}}=16$. All three models were trained with a batch size of 8 for 11k steps  [^4] In contrast to text-to-video, the video prediction task on the SkyTimelapse dataset has less diversity and converges faster. So we used smaller batch size and training steps..
Hyperparameters. For all the training, we used the DDPM (Ho et al., 2020) schedule with $T=1000$, $\beta_{1}=10^{-4}$, and $\beta_{T}=0.02$. The models were trained using AdamW (Loshchilov & Hutter, 2019) optimizer with a learning rate of 2e-5. At the inference stage, we used the improved DDPM schedule (Nichol & Dhariwal, 2021) with 100 steps. For text-to-video, we set the classifier-free guidance scale as 7.5.
@@ Appendix D Evaluation Details
Figure 11: Qualitative examples generated by GenLV (Wang et al., 2023a), StreamT2V (Henschel et al., 2025), OS-Fix, and our Ca2-VDM. We sampled 32 frames with an interval of 8 frames for display. Note that GenLV does not strictly follow the given first frame, since it was not finetuned on explicitly injected conditional frames. In the implementation of GenLV, we used DDIM inversion to build the initial noise based on the first frame.
@@ D.1 Datasets
MSR-VTT (Xu et al., 2016). we used its official test split which contains 2990 videos, with 20 manually annotated captions for each video. Following prior works (Ren et al., 2024; Zeng et al., 2024) and for fair comparisons, we randomly selected a caption for each video and generated 2990 videos for evaluation.
UCF101 (Soomro et al., 2012). As it only contains label names, we employed the descriptive text prompts from PYoCo (Ge et al., 2023), and generated 2048 samples with uniform distribution for each category following (He et al., 2022; Ge et al., 2023; Ren et al., 2024).
SkyTimelapse (Zhang et al., 2020). It is a time-lapse dataset showing dynamic sky scenes (e.g., cloudy sky with moving clouds). We used it for video prediction (i.e., without text input). Its training set contains 997 long timelapse videos, which are cut into 2392 short videos. Its test set contains 111 long timelapse videos, which are cut into 225 short videos. We trained the models on its training set and evaluated them on its test set.
@@ D.2 Quantitative Evaluation
Fréchet Video Distance (FVD) (Unterthiner et al., 2019) measures the similarity between generated and real videos based on the distributions on the feature space. We followed prior works (Blattmann et al., 2023b; Ge et al., 2022; Ren et al., 2024) to use a pretained I3D model [^5] https://github.com/songweige/TATS/blob/main/tats/fvd/i3d_pretrained_400.pt to extract the features. We used the codebase [^6] https://github.com/universome/stylegan-v from StyleGAN-V (Skorokhodov et al., 2021) to compute FVD statistics.
@@
Appendix E More Experiment Results
In Figure 11 and Figure 12, we show more qualitative examples from GenLV (Wang et al., 2023a), StreamT2V (Henschel et al., 2025), OS-Fix (Zheng et al., 2024), and Ca2-VDM. We can see that Ca2-VDM has comparable generation quality to existing SOTA models.
Figure 12: Qualitative examples from GenLV (Wang et al., 2023a), StreamT2V (Henschel et al., 2025), OS-Fix, and our Ca2-VDM. Yellow arrows highlight the consecutive frames having mutations.
In Table 6, we evaluated Ca2-VDM and OS-Ext on the VBench (Huang et al., 2024) benchmark. VBench is primarily designed for text-to-video evaluation. For our assessment, we selected four metrics: aesthetic quality, imaging quality, motion smoothness, and temporal flickering. The first two measure spatial (appearance) quality, and the last two assess temporal consistency. The results in Table 6 show that Ca2-VDM achieves comparable performance in both appearance quality and temporal consistency.
In Figure 13, we further compared the long-term content drift (i.e., error accumulation) between Ca2-VDM and the OS-Ext baseline. As a result, they show comparable visual quality. Both models exhibit a similar degree of error accumulation over time. Given our primary focus on efficiency, we conclude that Ca2-VDM matches the bidirectional baseline while being more efficient in both computation and storage for autoregressive video generation.
Table 6: VBench (Huang et al., 2024) evaluation on SkyTimelapse (Zhang et al., 2020) test set. The resolution of the generated video is $256\times 256$. Both models were evaluated with $P_{\text{max}}=25$ and $6$ autoregression steps.
Method Aesthetic Imaging Motion Temporal
Quality Quality Smoothness Flickering
OS-Ext 44.39 50.74 98.93 98.57
Ca2-VDM 44.30 50.55 97.59 97.14
Figure 13: Comparison between OS-Ext and Ca2-VDM in terms of long-term content drift (i.e., long-term quality degradation). Both models were trained on Sky-Timelapse (Zhang et al., 2020). Frame IDs are labeled at top-left corner.
Appendix F Limitations and Possible Future Directions
Causal Modeling in Pretraining. Currently, all the pretrained weights for video diffusion models (either UNet-based, e.g., ModelScore-T2V (Wang et al., 2023b), AnimateDiff (Guo et al., 2024b), or Transformer-based, e.g., Open-Sora (Zheng et al., 2024)) use bidirectional attention in their temporal modules. Our Ca2-VDM is built upon Open-Sora which was also pretrained using bidirectional attention. However, finetuning these bidirectionally pre-trained temporal modules using causal attention might be sub-optimal. The weights between bidirectional and causal temporal attention layers might have inherent gaps. Due to the limited computational resources, we did not conduct causal pretraining. Pretraining the VDM’s temporal modules from scratch (using causal attention) might have potential improvements.
Quality Degradation in Long-term Generation. As a common challenge, VDMs in long-term autoregressive generation suffer from frame appearance changes and quality degradation. Some works (Henschel et al., 2025; Zhang et al., 2023b) mitigate this issue by providing the VDM with the global appearance information extracted from the initial frame. However, during the long-term generation, video content may change and not all frames commit the same global appearance. In our setting, the long-term extendable context (i.e., early context from the KV-cache queue) helps mitigate the quality degradation, demonstrated by the results in Table 3 and Table 4. Further research on approaches addressing quality degradation is warranted and may hold potential significance for long-term video generation.